# LongVLM: Efficient Long Video Understanding via Large Language Models

### v2

@@ 4.1 Experimental Settings
Table 1: Comparison with state-of-the-art methods on video conversation benchmark [35] in terms of five evaluation aspects including Correctness of Information (CI), Detail Orientation (DO), Contextual Understanding (CU), Temporal Understanding (TU) and Consistency (C), along with the average scores across all evaluation metrics (Mean). We also report the dataset scale used for finetuning the model.
@@
LongVLM 100K 2.76 2.86 3.34 2.39 3.11 2.89
Table 2: Comparison with state-of-the-art methods on three zero-shot question answering datasets. We report the Accuracy (Acc.) and Score for the generated answer for each question. We also report the dataset scale used for finetuning the model.
@@
LongVLM 100K 47.6 3.3 59.8 3.3 70.0 3.8
Implementation details. We employ CLIP-ViT-L/14 [40] as the visual encoder and Vicuna-7B-v1.1 [7] as the LLM. We initialize them with the pretrained weights in LLaVA-7B-v1.1 [30]. We finetune the model on the Video-ChatGPT-100K instruction dataset [35] for 3 epochs, with a learning rate of $2\times e^{-5}$ and a batch size of 32. We only finetune the linear projection layer to align the visual features into the input space of the LLM, keeping both the visual encoder and LLM frozen. It takes three hours to train three epochs on 4 A100 80GB GPUs. During training and inference, we sample $T=100$ video frames for each video, and resize the frames to $224\times 224$ resolutions. We set $S=10$ for each video, and the number of tokens in each segment-level feature is $M=30$. We collect the $[CLS]$ tokens from the last five encoder layers and average them along the temporal axis, resulting in $E=5$ tokens as the global semantic features. Therefore, the length of visual tokens for a video sequence is $M\times S+E=305$.
@@ 4.3 Ablation Study
Effects of local feature aggregation. As discussed in Sec. 1, pooling operations or query aggregation might overlook local information in achieving fine-grained understanding in long-term videos. To this end, we introduce short-term segment-level features to retain local information and temporal structure within long-term videos. The first two rows in Tab. 3 present the effects for the design of using local features as the visual representations for videos. We compare the token merging module with local pooling operation using a pooling kernel and stride of [5,4,4] within each short-term segment. The proposed hierarchical merging module achieves higher scores over spatial-temporal pooling operation. This could be attributed to the dynamic aggregation mechanism via the token similarity in the merging module, while averaging pooling statically aggregates visual tokens within each small 3D window. Additionally, we observe that aggregating local features for short-term segments either improves or maintains comparable performance across all evaluation metrics compared to the SOTA models which extract global semantics only, highlighting the significance of preserving local features for short-term segments in long-term video understanding.
Table 3: Ablation of the local and global aggregation design. Pooling: Using local 3D pooling kernel to obtain local features; Merging: Using the proposed hierarchical token merging to obtain local features; $[{\bf L},{\bf G}]$: Concatenating local feature then global feature; $[{\bf G},{\bf L}]$: Concatenating global feature then local feature.
@@
E 1 5 10 15 20 24
CI 2.74 2.76 2.78 2.72 2.62 2.65
TU 2.32 2.39 2.30 2.16 2.28 2.22
C 3.04 3.11 3.04 2.97 2.93 2.81
Effects of global semantics integration. Inspired by the human visual system that using a combination of local and global information for recognizing video content [44], we propose to enhance visual representation by injecting global semantic features into local features. The last two rows in Tab. 3 demonstrate the effects of integrating global semantics. Compared to the first two rows, introducing global semantic features significantly enhances performance compared to models using local features only across all evaluation aspects. The notable improvements in Contextual Understanding (CU) and Consistency (C) underscore the significance of integrating global semantic information with local short-term features. Moreover, concatenating global features before local features yields better results than the opposite concatenation order. This arrangement allows each the local feature to access the global semantic information across the entire video by leveraging the causal attention mechanism in the LLM. Consequently, this enriches the contextual information of the local features and enhances the response consistency of the model.
@@
Figure 4: Two examples from Video-ChatGPT benchmark [35]. Text highlighted in bold green denotes correct content, while text in red indicates errors. The video durations are 42 seconds and 1 minute 39 seconds, respectively.
@@
Effects of $M$. We report the model performance on VideoChatGPT benchmark and ANET-QA task on the selection of $M$, i.e., $M=\{10,20,30,40\}$, keeping the same number of global semantic tokens in Tab. 4. For ANET-QA task, we also report the averaging GPU memory usage for generating each response. In general, the token length involves a trade-off between memory costs and performance. A shorter token sequence reduces computational costs for generating a single new token using LLM, thereby lowering memory costs for generating responses to individual user queries. However, it may also lead to insufficient visual information for generating accurate responses. The performance of our model is beneficial from the suitable length of visual tokens. Specifically, using $M=10$ leads to the lowest performance in both tasks, while it still shows a comparable performance compared to the SOTA models, indicating the effectiveness of our model design. Increasing $M$ from 10 to 40 results in a significant improvement in terms of most evaluation aspects, while the setting of $M=40$ leads to neglecting improvement but requires more memory cost compared to $M=30$. Therefore, we choose $M=30$ for our model.
@@ 5 Conclusion
In this work, we have introduced LongVLM, an effective and efficient multimodal LLM designed for long-term video understanding. By extracting local features for short-term segments, we efficiently model local dependencies while preserving the temporal structure of sequential events in long-term video sequences. Through the integration of local and global information, LongVLM captures detailed information and provides consistent and accurate responses for long-term videos.

### v3

@@ 4.1 Experimental Settings
Table 1: Comparison with state-of-the-art methods on video conversation benchmark [39] in terms of five evaluation aspects and the average scores across all aspects (Mean). We also report the dataset scale used for finetuning the model.
@@
Ours 100K 2.76 2.86 3.34 2.39 3.11 2.89
Table 2: Comparison with state-of-the-art methods on three zero-shot question answering datasets. We report the Accuracy (Acc.) and Score for the generated answer for each question, and the dataset scale used for finetuning the model.
@@
Ours 100K 47.6 3.3 59.8 3.3 70.0 3.8
Implementation details. We employ CLIP-ViT-L/14 [44] as the visual encoder and Vicuna-7B-v1.1 [8] as the LLM. We initialize them with the pretrained weights in LLaVA-7B-v1.1 [34]. We finetune the model on the Video-ChatGPT-100K instruction dataset [39] for 3 epochs, with a learning rate of $2e^{-5}$ and a batch size of 32. We only finetune the linear projection layer to align the visual features into the input space of the LLM, keeping both the visual encoder and LLM frozen. It takes three hours to train three epochs on 4 A100 80GB GPUs. During training and inference, we sample $T=100$ video frames for each video, and resize the frames to $224\times 224$ resolutions. We set $C=16$, the same to the number of heads of CLIP-ViT/L-14. We set $S=10$ for each video, and the number of tokens in each segment-level feature is $M=30$. We collect the [CLS] tokens from the last five encoder layers and average them along the temporal axis, resulting in $E=5$ tokens as the global semantic features. Therefore, the length of visual tokens for a video sequence is $M\times S+E=305$.
@@ 4.3 Ablation Study
Effects of local feature aggregation. As discussed in Sec. 1, pooling operations or query aggregation might overlook local information in achieving fine-grained understanding in long-term videos. To this end, we introduce short-term segment-level features to retain local information and temporal structure within long-term videos. The first two rows in Tab. 3 present the effects for the design of using local features as the visual representations for videos. We compare the token merging module with a local pooling operation. Specifically, we apply a 3D average pooling operation within each short-term segment, using a kernel size and stride of (5, 4, 4) to the temporal, height, and width dimensions, respectively. The proposed hierarchical merging module achieves higher scores over spatial-temporal pooling operation. This could be attributed to the dynamic aggregation mechanism via the token similarity in the merging module, while averaging pooling statically aggregates visual tokens within each small 3D window. Additionally, we observe that aggregating local features for short-term segments either improves or maintains comparable performance across all evaluation metrics compared to the SOTA models which extract global semantics only, highlighting the significance of preserving local features for short-term segments in long-term video understanding.
Table 3: Ablation of the local and global aggregation design. Pooling: Using 3D average pooling operation to obtain local features; Merging: Using the proposed hierarchical token merging to obtain local features; $[{\bf L},{\bf G}]$: Concatenating local feature then global feature; $[{\bf G},{\bf L}]$: Concatenating global feature then local feature.
@@
E CI DO CU TU C Mean
1 2.74 2.85 3.23 2.32 3.04 2.83
5 2.76 2.86 3.34 2.39 3.11 2.89
10 2.78 2.86 3.24 2.30 3.04 2.84
15 2.72 2.82 3.16 2.16 2.97 2.77
20 2.63 2.77 3.11 2.28 2.93 2.74
24 2.65 2.75 3.08 2.22 2.81 2.70
@@ E CI DO CU TU C Mean
Figure 4: Two examples from Video-ChatGPT benchmark [39]. Text highlighted in bold green denotes correct content, while text in red indicates errors.
Effects of global semantics integration. Inspired by the human visual system that using a combination of local and global information for recognizing video content [49], we propose to enhance visual representation by injecting global semantic features into local features. The last two rows in Tab. 3 demonstrate the effects of integrating global semantics. Compared to the first two rows, introducing global semantic features significantly enhances performance compared to models using local features only across all evaluation aspects. The notable improvements in Contextual Understanding (CU) and Consistency (C) underscore the significance of integrating global semantic information with local short-term features. Moreover, concatenating global features before local features yields better results than the opposite concatenation order. This arrangement allows each the local feature to access the global semantic information across the entire video by leveraging the causal attention mechanism in the LLM. Consequently, this design enriches the contextual information of the local features and enhances the response consistency of the model.
@@
Effects of $M$. We report the model performance on VideoChatGPT benchmark and ANET-QA task on the selection of $M$, i.e., $M=\{10,20,30,40\}$, keeping the same number of global semantic tokens in Tab. 4. For ANET-QA, we also report the averaging GPU memory usage for generating each answer. In general, the token length involves a trade-off between memory costs and performance. A shorter token sequence reduces computational costs for generating a single new token using LLM, thereby lowering memory costs for generating responses to individual user queries. However, it may also lead to insufficient visual information for generating accurate responses. The performance of our model is beneficial from the suitable length of visual tokens. Increasing $M$ from 10 to 40 results in a significant improvement in terms of most evaluation aspects, while the setting of $M=40$ leads to neglecting improvement but requires more memory cost compared to $M=30$. Therefore, we choose $M=30$ for our model.
@@ 5 Conclusion
In this work, we have introduced LongVLM, an effective and efficient VideoLLM designed for long-term video understanding. By extracting local features for short-term segments, we efficiently model local dependencies while preserving the temporal structure of sequential events in long-term video sequences. Through the integration of local and global information, LongVLM captures detailed information and provides consistent and accurate responses for long-term videos.
@@
Appendix
We organize our supplementary material as follows.
  • In Sec. 6, we provide additional ablation results.
  • In Sec. 7, we provide more generation results from the proposed LongVLM.
6 Additional Results
Effects of temperature. We show the effects of temperature settings in the LLM on both performance scores and averaged word counts in responses on Video-ChatGPT benchmark [39]. As demonstrated in Tab. 6 and Fig. 6, lower temperatures lead to shorter but more accurate responses, with tokens that are more deterministic and closely aligned with the most relevant predictions. However, increasing the temperature from 0.2 to 2.0 lowers performance and raises word counts, leading to longer, more creative responses that may be irrelevant and inaccurate. Therefore, we set the temperature at 0.2 in our model.
Table 6: Effects of temperature. We report the Correctness of Information (CI, Temporal Understanding (TU), and Consistency (C) by varying temperature in {0.01, 0.1, 0.2, 0.5, 1.0, 2.0}.
Temperature 0.01 0.1 0.2 0.5 1.0 2.0
CI 2.74 2.76 2.76 2.69 2.35 1.21
TU 2.22 2.16 2.39 2.20 1.99 1.25
C 2.85 2.89 3.11 2.62 2.21 1.01
(a) CI
(b) TU
(c) C
Figure 6: Average word counts in responses for various temperature settings. "Q1" and "Q2" denote two questions addressing similar perspectives in the Consistency (C) evaluation metrics.
More results on zero-shot QA. We evaluate our model in a zero-shot manner on subsets of two datasets: (1) Egoschema [40], a long-form video QA dataset derived from Ego4D [19]; and (2) MAD [46], an audio description (AD) generation benchmark that requires understanding long contexts in hour-long movies. We sample videos at 1 fps for Egoschema and 5 fps for MAD. We report the accuracy (%) for Egoschema and the response score (ranging from 0 to 5) for MAD in Tab. 7, demonstrating the effectiveness of the proposed LongVLM on longer video datasets.
Table 7: Model performance on Egoschema and MAD.
Method Egoschema MAD
VideoChat 46.6 1.90
Video LLaMA 38.8 1.86
Video-ChatGPT 49.6 1.93
BT-Adapter 54.6 2.14
LongVLM 57.6 2.21
Additional ablation results on zero-shot QA. We verify the local feature aggregation and global semantics integration strategies on ANET-QA [64] and Egoschema [40]. We report the accuracy (%) in Tab. 8. The results demonstrate the effectiveness of token merging module compared to the pooling operation on longer videos. Additionally, concatenating global features before local features yields better results than the opposite concatenation order.
Table 8: Ablation of the local and global aggregation design.
Variants ANET-QA Egoschema
Pooling 45.4 48.0
Merging 46.5 52.4
[L, G] 47.0 56.0
[G, L] 47.6 57.6
7 Additional Qualitative Examples
We provide more generated examples from VideoChatGPT benchmark, ANET-QA and MSRVTT-QA. The proposed LongVLM can generate precise and accurate responses that correspond to video content and correctly answer questions about detailed information in videos.